Fermi National Laboratory
 

Accelerator Update for November 19, 2007 - November 21, 2007

 

Monday November 19

The day shift began with Operations monitoring the Tevatron (TeV) store 5731 (34.02E30) and the Recycler stash (230.6E10), with the Antiproton Source (Pbar) stacking (42.4E10), with the Main Injector (MI), Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams, with NuMI off due to a kicker problem, and with the SciBooNE / MiniBooNE experiments taking beam.

At 8:12 AM, experts reported that the NuMI kicker (KPS6N) had an oil leak that must be repaired. The repair might take 16 to 24 hours to complete.

Operations terminated store 5731 at 10:01 AM.

Operations reported at 10:16 AM that cryo system technicians were replacing a card in the TeV sector A3 thermometry crate.

At 10:19 AM, the MI power supply interlock loop tripped off on a ground fault indication. It reset.

TeV experts began a study at 12:16 PM. The study ended about 50 minutes later.

Operations put the TeV into shot setup at 1:16 PM.

At 3:22 PM, the TeV suffered a sector F11 and A11 quench at the start of the squeeze. Experts investigated.

Operations put the TeV through a dry squeeze at 11:35 PM.

NuMI experts reported at 11:45 PM that they found the oil leak on their extraction kicker. The leak was on a nut that had loosened due to arcing and ruined its oil seal.

  

Operations reported at midnight that the Recycler had stashed antiprotons twice during the evening shift.

Tuesday November 20

The midnight shift began with Operations monitoring the stash (176.8E10), with Pbar stacking (56.0E10), and with Sci /MiniBooNE taking beam.

At 12:19 AM, Operations reported that a TeV power supply (T:VE12) had tripped off due to an over temperature indication and would not reset. Operators replaced its regulator, but it still didn’t reset. An expert was called in.

A TeV power supply expert reported at 2 AM that he had found and repaired a clogged LCW supply hose on VE12.

At 2:10 AM, Operations reported that the TeV sectors A3 and A4 thermometry crates were not responding and would not reset. An expert was called in.

At 2:44 AM, the Pelletron thermocouple came into alarm. The temperature for the Pelletron is above its limit, but not in danger of tripping off. Operations notified experts.

Operations reported at 2:44 AM that they were able to reset the A3 and A4 thermometry crates.

Operations put the TeV through a dry squeeze at 2:58 AM.

At 3:01 AM, Operations reported that the A3 and A4 thermometry crates tripped off.

Operations reported at 3:16 AM that a cryo system expert swapped out the hub module for the sector A thermometry crates and the power supply for the A3 crate.

Operations put the TeV through a proton only wet squeeze at 4:03 AM.

Operations put the TeV into shot setup at 4:34 AM.

Operations established store 5737 at 6:28 AM with an initial luminosity of 173.36E30.

Operations reported at 7:57 AM that the Recycler had stashed three times during the midnight shift.

Operations reported at 10:01 AM that experts had completed the MI DCCT testing.

Instrumentation experts began working on the MI BPMs at 11:03 AM.

Pbar experts began a reverse proton study at 11:59 AM.

MI experts began various studies at 12:47 PM.

Operations reported at 1:08 PM that the NuMI kicker (KPS6N) was running at 25 kV.

Booster experts began a study around 1:57 PM.

NuMI kicker experts reported that their work on KPS6N was finished. Operations tried to send beam to NuMI at 3:53 PM, but MIRF10 kept tripping off.

Operations reported at 4:59 PM that the Recycler had stashed twice during the day shift.

A RF expert investigated the MIRF10 problem and had it back on line by 4:29 PM.

TeV experts began separator scans at 4:31 PM.

At 4:55 PM, MIRF10 tripped off. An expert investigated.

At 5:02 PM, Operations reported that MIRF10 tripped a yard breaker. The Duty Electrician was sent out to reset the breaker.

Operations reported at 5:18 PM that TeV experts had completed their separator scans.

Operations established beam to NuMI at 5:42 PM.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Wednesday November 21

The midnight shift began with Operations monitoring store 5737 (40.03E30) and a stash (144.8E10), with Pbar stacking (48.4E10), and with NuMI and Sci/MiniBooNE taking beam.

At 1:03 AM, MI tripped off due to a MI-50 quadrupole magnet bus to ground indication. It reset.

At 3:45 AM, the NuMI beam permit tripped off due to a number of quadrupole magnet power supplies not ramping. Operations called in an expert.

The power supply expert replaced a NuMI regulator 15-volt power supply. NuMI resumed taking beam at 4:50 AM.

At 5:03 AM, the NuMI beam permit tripped off due to an absorber LCW water skid tripping off. Operations called in an expert.

Operations terminated store 5737 at 8 AM.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday and the long Thanksgiving Weekend

The plans for today and the weekend are to stack, stash, and store. CDF reported that they have a problem with their COT, which will take 8-12 hours to repair. The Run Coordinator may allow CDF to access their collision hall and make repairs on Friday, 11/23/07. If this occurs, he will also allow other machines a chance to access and perform maintenance.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/21/2007 -->    email Fermilab